# Finite Sample Analysis of Distributional TD Learning with Linear Function Approximation

### v1

@@ C.1 Linear-Categorical Parametrization is an Isometry
By Proposition B.1,
$\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{w}}_{1}},\bm{\eta}_{{\bm{w}}_{2}}}\right)=$ $\iota_{K}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{\left\|{{\bm{C}}\left({{\bm{p}}_{{\bm{w}}_{1}}(s)-{\bm{p}}_{{\bm{w}}_{2}}(s)}\right)}\right\|^{2}}\right]$
$=$ $\iota_{K}\operatorname{tr}\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{\frac{1}{2}}\left({{\bm{W}}_{1}-{\bm{W}}_{2}}\right){\bm{C}}^{\intercal}{\bm{C}}\left({{\bm{W}}_{1}-{\bm{W}}_{2}}\right)^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{\frac{1}{2}}}\right)$
$=$ $\iota_{K}\left\|{{\bm{w}}_{1}-{\bm{w}}_{2}}\right\|^{2}_{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}.$
C.2 Linear-Categorical Projection Operator
Proposition 3.1 is an immediate corollary of the following lemma.
Lemma C.1.
For any $\bm{\eta}\in\left({{\mathscr{P}}^{\mathsf{sign}}}\right)^{{\mathcal{S}}}$, ${\bm{w}}\in{\mathbb{R}}^{dK}$ and $s\in{\mathcal{S}}$, it holds that
$\nabla_{{\bm{W}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}}(s),{\eta}(s)}\right)$ $=2\iota_{K}{\bm{\phi}}(s)\left({{\bm{p}}_{{\bm{w}}}(s)-{\bm{p}}_{\bm{\eta}}(s)}\right)^{\intercal}{\bm{C}}^{\intercal}{\bm{C}}$
$=2\iota_{K}{\bm{\phi}}(s)\left[{{\bm{\phi}}(s)^{\intercal}{\bm{W}}+\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)^{\intercal}}\right]{\bm{C}}^{\intercal}{\bm{C}}.$
@@ C.2 Linear-Categorical Projection Operator
$\nabla_{{\bm{W}}}\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{w}}},\bm{\eta}}\right)$ $={\mathbb{E}}_{s\sim\mu_{\pi}}\left[{\nabla_{{\bm{W}}}\ell_{2}^{2}\left({{\eta}_{{\bm{w}}}(s),{\eta}(s)}\right)}\right]$
$=2\iota_{K}\left[{{\bm{\Sigma}}_{{\bm{\phi}}}{\bm{W}}+{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)^{\intercal}}\right]}\right]{\bm{C}}^{\intercal}{\bm{C}}.$
@@
According to Proposition B.2, one has
$\ell_{2}^{2}\left({{\eta}_{{\bm{w}}}(s),{\eta}(s)}\right)=\ell_{2}^{2}\left({{\eta}_{{\bm{w}}}(s),{\bm{\Pi}}_{K}{\eta}(s)}\right)+\ell_{2}^{2}\left({{\bm{\Pi}}_{K}{\eta}(s),{\eta}(s)}\right).$

### v2

@@ D.1 Linear-Categorical Parametrization is an Isometry
For any $\bm{\eta}_{{\bm{\theta}}}\in{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}$, we denote ${\bm{F}}_{{\bm{\theta}}}(s)=\left({F_{k}(s;{\bm{\theta}})}\right)_{k=0}^{K-1}\in{\mathbb{R}}^{K}$, then it holds that
$\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{\theta}}_{1}},\bm{\eta}_{{\bm{\theta}}_{2}}}\right)=$ $\iota_{K}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{\left\|{{{\bm{F}}_{{\bm{\theta}}_{1}}(s)-{\bm{F}}_{{\bm{\theta}}_{2}}(s)}}\right\|^{2}}\right]$
$=$ $\iota_{K}\operatorname{tr}\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{\frac{1}{2}}\left({{\bm{\Theta}}_{1}-{\bm{\Theta}}_{2}}\right)\left({{\bm{\Theta}}_{1}-{\bm{\Theta}}_{2}}\right)^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{\frac{1}{2}}}\right)$
$=$ $\iota_{K}\left\|{{\bm{\theta}}_{1}-{\bm{\theta}}_{2}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}.$
D.2 Linear-Categorical Projection Operator
Proposition 3.1 is an immediate corollary of the following lemma. For any $\nu\in{\mathscr{P}}^{\mathsf{sign}}_{K}$, we define ${\bm{F}}_{\nu}=\left({F_{k}\left({\nu}\right)}\right)_{k=0}^{K-1}=\left({\nu\left({[0,x_{k}]}\right)}\right)_{k=0}^{K-1}\in{\mathbb{R}}^{K}$, and for any $\bm{\eta}\in\left({{\mathscr{P}}^{\mathsf{sign}}}\right)^{{\mathcal{S}}}$, we define ${\bm{p}}_{\bm{\eta}}(s)={\bm{p}}_{\Pi_{K}\eta(s)}$ and ${\bm{F}}_{\bm{\eta}}(s)={\bm{F}}_{\Pi_{K}\eta(s)}$.
Lemma D.1.
For any $\bm{\eta}\in\left({{\mathscr{P}}^{\mathsf{sign}}}\right)^{{\mathcal{S}}}$, ${\bm{\theta}}\in{\mathbb{R}}^{dK}$ and $s\in{\mathcal{S}}$, it holds that
$\nabla_{{\bm{\Theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{\theta}}}(s),{\eta}(s)}\right)$ $=2\iota_{K}{\bm{\phi}}(s)\left({{\bm{F}}_{{\bm{\theta}}}(s)-{\bm{F}}_{\bm{\eta}}(s)}\right)^{\top}$ (19)
$=2\iota_{K}{\bm{\phi}}(s)\left[{{\bm{\phi}}(s)^{\top}{\bm{\Theta}}+\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)^{\top}{\bm{C}}^{\top}}\right].$
@@ D.2 Linear-Categorical Projection Operator
$\nabla_{{\bm{\Theta}}}\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{\theta}}},\bm{\eta}}\right)$ $={\mathbb{E}}_{s\sim\mu_{\pi}}\left[{\nabla_{{\bm{\Theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{\theta}}}(s),{\eta}(s)}\right)}\right]$
$=2\iota_{K}\left[{{\bm{\Sigma}}_{{\bm{\phi}}}{\bm{\Theta}}+{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)^{\top}}\right]{\bm{C}}^{\top}}\right].$
@@
According to Proposition C.2, one has
$\ell_{2}^{2}\left({{\eta}_{{\bm{\theta}}}(s),{\eta}(s)}\right)=\ell_{2}^{2}\left({{\eta}_{{\bm{\theta}}}(s),{\bm{\Pi}}_{K}{\eta}(s)}\right)+\ell_{2}^{2}\left({{\bm{\Pi}}_{K}{\eta}(s),{\eta}(s)}\right).$
@@
$\nabla_{{\bm{\theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{\theta}}}(s),{\eta}(s)}\right)$ $=\nabla_{{\bm{\theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{\theta}}}(s),{\bm{\Pi}}_{K}{\eta}(s)}\right)$
$=\iota_{K}\nabla_{{\bm{\theta}}}\left\|{{{\bm{F}}_{{\bm{\theta}}}(s)-{\bm{F}}_{\bm{\eta}}(s)}}\right\|^{2}$
$=2\iota_{K}\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right)\left({\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s)}\right)^{\top}{\bm{\theta}}+{\bm{C}}\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)}\right)$
$=2\iota_{K}\left\{{\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right)}\right]{\bm{\theta}}+\left[{\left({{\bm{C}}\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)}\right)\otimes{\bm{\phi}}(s)}\right]}\right\}.$
We also have the following matrix representation:
$\nabla_{{\bm{\Theta}}}\ell_{2}^{2}\left({{\eta}_{{\bm{\theta}}}(s),{\eta}(s)}\right)$ $=2\iota_{K}{\bm{\phi}}(s)\left({{\bm{F}}_{{\bm{\theta}}}(s)-{\bm{F}}_{\bm{\eta}}(s)}\right)^{\top}$
$=2\iota_{K}{\bm{\phi}}(s)\left[{{\bm{\phi}}(s)^{\top}{\bm{\Theta}}+\left({\frac{1}{K+1}\bm{1}_{K}-{\bm{p}}_{\bm{\eta}}(s)}\right)^{\top}{\bm{C}}^{\top}}\right].$
Proposition D.2.
For any $\bm{\eta}\in\left({{\mathscr{P}}^{\mathsf{sign}}}\right)^{{\mathcal{S}}}$ and $\bm{\eta}_{{\bm{\theta}}}\in{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}$, it holds that
$\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta},\bm{\eta}_{{\bm{\theta}}}}\right)=\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta},{\bm{\Pi}}_{K}\bm{\eta}}\right)+\ell_{2,\mu_{\pi}}^{2}\left({{\bm{\Pi}}_{K}\bm{\eta},{\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}}\right)+\ell_{2,\mu_{\pi}}^{2}\left({{\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta},\bm{\eta}_{{\bm{\theta}}}}\right).$
The proof is straightforward and almost the same as that of Proposition C.2 if we utilize the affine structure.
D.3 Linear-Categorical Projected Bellman Equation
To derive the result, the following proposition characterizing ${\bm{\Pi}}_{K}\bm{\mathcal{T}}^{\pi}\bm{\eta}_{{\bm{\theta}}}$ is useful, whose proof can be found in Appendix D.4.
Proposition D.3.
For any ${\bm{\theta}}\in{\mathbb{R}}^{dK}$ and $s\in{\mathcal{S}}$, we abbreviate ${\bm{p}}_{\bm{\mathcal{T}}^{\pi}\bm{\eta}_{\bm{\theta}}}(s)$ as $\tilde{{\bm{p}}}_{{\bm{\theta}}}(s)$, then
$\tilde{{\bm{p}}}_{{\bm{\theta}}}(s)=$ $\left({\tilde{p}_{k}(s;{\bm{\theta}})}\right)_{k=0}^{K-1}={\mathbb{E}}\left[{\tilde{{\bm{G}}}(r_{0}){\bm{C}}^{-1}{\bm{\Theta}}^{\top}{\bm{\phi}}(s_{1})\Big{|}s_{0}=s}\right]+\frac{1}{K+1}\sum_{j=0}^{K}{\mathbb{E}}\left[{{\bm{g}}_{j}(r_{0})\Big{|}s_{0}=s}\right].$
Combining this proposition with Proposition 3.1, we know that ${\bm{\Theta}}^{\star}$ is the unique solution to the following system of linear equations for ${\bm{\Theta}}\in{\mathbb{R}}^{d\times K}$
${\bm{\Theta}}=$ ${\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({\tilde{\bm{p}}_{{\bm{\theta}}}(s)-\frac{1}{K+1}\bm{1}_{K}}\right)^{\top}{\bm{C}}^{\top}}\right]$
$=$ ${\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({\frac{1}{K+1}\bm{1}_{K}-{\mathbb{E}}\left[{\tilde{{\bm{G}}}(r_{0}){\bm{C}}^{-1}{\bm{\Theta}}^{\top}{\bm{\phi}}(s_{1})\Big{|}s_{0}=s}\right]-\frac{1}{K+1}\sum_{j=0}^{K}{\mathbb{E}}\left[{{\bm{g}}_{j}(r_{0})\Big{|}s_{0}=s}\right]}\right)^{\top}{\bm{C}}^{\top}}\right]$
$=$ ${\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Theta}}\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)^{\top}}\right]+\frac{1}{K+1}{\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r)-\bm{1}_{K}}\right)^{\top}{\bm{C}}^{\top}}\right],$
or equivalently,
${\bm{\Sigma}}_{{\bm{\phi}}}{\bm{\Theta}}-{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Theta}}\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)^{\top}}\right]=\frac{1}{K+1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r)-\bm{1}_{K}}\right)^{\top}{\bm{C}}^{\top}}\right],$
which is the desired conclusion. The uniqueness and existence of the solution is guaranteed by the fact that the LHS is an invertible linear transformation of ${\bm{\Theta}}$, which is justified by Eqn. (28).
D.4 Proof of Proposition D.3
Proof.
$\tilde{p}_{k}(s;{\bm{\theta}})=$ $p_{k}\left({\left[{\bm{\mathcal{T}}^{\pi}\bm{\eta}_{\bm{\theta}}}\right](s)}\right)$ (20)
$=$ ${\mathbb{E}}_{X\sim\left[{\bm{\mathcal{T}}^{\pi}\bm{\eta}_{\bm{\theta}}}\right](s)}\left[{\left({1-\left|{\frac{X-x_{k}}{\iota_{K}}}\right|}\right)_{+}}\right]$
$=$ ${\mathbb{E}}\left[{{\mathbb{E}}_{G\sim\eta_{{\bm{\theta}}}(s_{1})}\left[{\left({1-\left|{\frac{r_{0}+\gamma G-x_{k}}{\iota_{K}}}\right|}\right)_{+}}\right]\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{\sum_{j=0}^{K}p_{j}(s_{1};{\bm{\theta}})\left({1-\left|{\frac{r_{0}+\gamma x_{j}-x_{k}}{\iota_{K}}}\right|}\right)_{+}\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{\sum_{j=0}^{K}p_{j}(s_{1};{\bm{\theta}})g_{j,k}(r_{0})\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{g_{K,k}(r_{0})+\sum_{j=0}^{K-1}p_{j}(s_{1};{\bm{\theta}})\left({g_{j,k}(r_{0})-g_{K,k}(r_{0})}\right)\Big{|}s_{0}=s}\right].$
Hence, let ${\bm{W}}={\bm{\Theta}}{\bm{C}}^{-\top}$ and ${\bm{w}}=\mathsf{vec}({\bm{W}})=\left({{\bm{C}}^{-1}\otimes{\bm{I}}_{d}}\right){\bm{\theta}}$ (see Appendix D.6 for their meaning), then
$\tilde{{\bm{p}}}_{{\bm{\theta}}}(s)=$ $\left({\tilde{p}_{k}(s;{\bm{\theta}})}\right)_{k=0}^{K-1}$
$=$ ${\mathbb{E}}\left[{\begin{bmatrix}g_{K,1}(r_{0})\\ \vdots\\ g_{K,{K-1}}(r_{0})\end{bmatrix}+\sum_{j=0}^{K-1}p_{j}(s_{1};{\bm{\theta}})\begin{bmatrix}g_{j,1}(r_{0})-g_{K,1}(r_{0})\\ \vdots\\ g_{j,K-1}(r_{0})-g_{K,K-1}(r_{0})\end{bmatrix}\Bigg{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{{\bm{g}}_{K}(r_{0})+\sum_{j=0}^{K-1}p_{j}(s_{1};{\bm{\theta}})\left({{\bm{g}}_{j}(r_{0})-{\bm{g}}_{K}(r_{0})}\right)\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{{\bm{g}}_{K}(r_{0})+\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\top}\otimes{\bm{g}}_{K}(r_{0})}\right){\bm{p}}_{{\bm{\theta}}}(s_{1})\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{{\bm{g}}_{K}(r_{0})+\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\top}\otimes{\bm{g}}_{K}(r_{0})}\right)\left[{\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s_{1})}\right)^{\top}{\bm{w}}+\frac{1}{K+1}\bm{1}_{K}}\right]\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\top}\otimes{\bm{g}}_{K}(r_{0})}\right)\left({{\bm{I}}_{K}\otimes{\bm{\phi}}(s_{1})}\right)^{\top}\Big{|}s_{0}=s}\right]{\bm{w}}$
$\quad+{\mathbb{E}}\left[{{\bm{g}}_{K}(r_{0})+\frac{1}{K+1}\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\top}\otimes{\bm{g}}_{K}(r_{0})}\right)\bm{1}_{K}\Big{|}s_{0}=s}\right]$
$=$ ${\mathbb{E}}\left[{\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\top}\otimes{\bm{g}}_{K}(r_{0})}\right)\otimes{\bm{\phi}}(s_{1})^{\top}\Big{|}s_{0}=s}\right]{\bm{w}}+\frac{1}{K+1}\sum_{j=0}^{K}{\mathbb{E}}\left[{{\bm{g}}_{j}(r_{0})\Big{|}s_{0}=s}\right],$
or equivalently,
$\tilde{{\bm{p}}}_{{\bm{\theta}}}(s)$ $={\mathbb{E}}\left[{\tilde{{\bm{G}}}(r_{0}){\bm{W}}^{\top}{\bm{\phi}}(s_{1})\Big{|}s_{0}=s}\right]+\frac{1}{K+1}\sum_{j=0}^{K}{\mathbb{E}}\left[{{\bm{g}}_{j}(r_{0})\Big{|}s_{0}=s}\right]$
$={\mathbb{E}}\left[{\tilde{{\bm{G}}}(r_{0}){\bm{C}}^{-1}{\bm{\Theta}}^{\top}{\bm{\phi}}(s_{1})\Big{|}s_{0}=s}\right]+\frac{1}{K+1}\sum_{j=0}^{K}{\mathbb{E}}\left[{{\bm{g}}_{j}(r_{0})\Big{|}s_{0}=s}\right].$
D.5 Proof of Proposition 3.2
Proof.
By the basic inequality (Lemma G.1), we only need to show
$\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}^{\pi},\bm{\eta}_{{\bm{\theta}}^{\star}}}\right)\leq$ $\frac{\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}^{\pi},{\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}^{\pi}}\right)}{1-\gamma}$
where we used Bellemare et al. [4, Proposition 9.18 and Eqn. (5.28)]. ∎
D.6 Linear-Categorical Parametrization with Probability Mass Function Representation
We introduce new notations for linear-categorical parametrization with probability mass function (PMF) representation. Let ${\bm{W}}:={\bm{\Theta}}{\bm{C}}^{-\top}=\left({{\bm{\theta}}(0),{\bm{\theta}}(1)-{\bm{\theta}}(0),\cdots,{\bm{\theta}}(K-1)-{\bm{\theta}}(K-2)}\right)\in{\mathbb{R}}^{d\times K}$ and the vectorization of ${\bm{W}}$, ${\bm{w}}:=\mathsf{vec}\left({{\bm{W}}}\right)=\left({{\bm{C}}^{-1}\otimes{\bm{I}}_{d}}\right){\bm{\theta}}\in{\mathbb{R}}^{dK}$. We abbreviate ${\bm{p}}_{\bm{\eta}_{\bm{\theta}}}$ as ${\bm{p}}_{{\bm{w}}}$ in this section. Then by Lemma A.2, for any $s\in{\mathcal{S}}$, it holds that
${\bm{p}}_{{\bm{w}}}(s)={\bm{W}}^{\top}{\bm{\phi}}(s)+(K+1)^{-1}\bm{1}_{K}.$ (21)
PMF and CDF representations are equivalent because ${\bm{C}}$ is invertible.
For any $\bm{\eta}_{{\bm{w}}_{1}},\bm{\eta}_{{\bm{w}}_{2}}\in{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}$, by Proposition C.1,
$=$ $\iota_{K}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{\left\|{{\bm{C}}\left({{\bm{p}}_{{\bm{w}}_{1}}(s)-{\bm{p}}_{{\bm{w}}_{2}}(s)}\right)}\right\|^{2}}\right]$
$=$ $\iota_{K}\operatorname{tr}\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{\frac{1}{2}}\left({{\bm{W}}_{1}-{\bm{W}}_{2}}\right){\bm{C}}^{\top}{\bm{C}}\left({{\bm{W}}_{1}-{\bm{W}}_{2}}\right)^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{\frac{1}{2}}}\right)$
$=$ $\iota_{K}\left\|{{\bm{w}}_{1}-{\bm{w}}_{2}}\right\|^{2}_{\left({{\bm{C}}^{\top}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}},$
hence the affine space $\left({{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K},\ell_{2,\mu_{\pi}}}\right)$ is isometric with the Euclidean space $\left({{\mathbb{R}}^{Kd},\sqrt{\iota_{K}}\left\|{\cdot}\right\|_{\left({{\bm{C}}^{\top}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}}\right)$ if we consider the PMF representation.
Following the proof of Lemma D.1 in Appendix D.2, we can also derive the gradient when we use the PMF parametrization: